Gomi Waste Disposal is evaluating a project that would require an initial investment of 56,100 dollars today. The project is then expected to produce annual cash flows that grow by 3.9 percent per year forever. The first annual cash flow is expected in 1 year and is expected to be 2,290 dollars. The project’s internal rate of return is 7.98 percent and its cost of capital is 10.73 percent. What is the net present value (NPV) of the project?
Answer #1

For calculating NPV, we have to use cost of capital and not IRR.

Year 1 cashflow (CF1) = $ 2,290

Annual growth rate (g) = 3.9%

Cost of Capital (K) = 10.73%

Using Gordon's Dividend Growth model,

PV of inflows = CF1 / (K-g)

PV of inflows = 2290 / (10.73%-3.90%)

PV of inflows = $ 33,528.55

Net Present Value = PV of inflows - PV of outflows

Net Present Value = 33,528.55 - 56,100

Net Present Value (NPV) = (-$ 22,571.45)

Thumbs up please if satisfied. :)

Comment if further doubts in above.